# Chatbots as social companions: How people perceive consciousness, human likeness, and social health benefits in machines

### v4

@@ 3.1 Perceptions of the human-chatbot relationship
Figure 1: Distributions of responses to questions pertaining to social health measures for companion chatbot users and non-users. Participants were asked to rate how a relationship with a companion chatbot harmed or helped (user group) or might harm or help (non-user group) their social interactions, relationships with family and friends, and self-esteem. The x-axis shows the Likert-scale response options from 1-7 (“Very harmful” to “Very helpful”), and the y-axis shows the frequency of responses.
@@
Figure 2: Distributions of responses pertaining to hypothetical changes to the companion chatbot or to dependence on the companion chatbot. Participants were asked to rate their perceptions of if the companion chatbot really had emotions, if it developed into a living being, or if they depended on it a lot. The x-axis shows the Likert-scale response options from 1-7, and the y-axis shows the frequency of responses.
@@
We assessed how users and non-users perceived the impact of a chatbot relationship on their social health. Figure 1A shows the results for companion chatbot users, when asked whether they judged their relationship with the chatbot to be harmful or helpful for their general social interactions with other people. The distribution is skewed toward the “Very helpful” side of the 7-point scale (M=5.16), and no one chose the lowest rating. Figure 1B shows the results for companion chatbot users on whether they judged their relationship with the chatbot to be harmful or helpful specifically for their relationships with their families and friends. Again, the distribution is skewed toward the positive side (M=4.84). The mode of responses in Figure 1A and 1B correspond to “Neutral,” but the means of these distributions correspond to scores between “Slightly helpful” and “Moderately helpful.” Figure 1C shows the results for how companion chatbot users perceived the chatbot relationship to have impacted their self-esteem, and the distribution is even more skewed to the positive side (M=5.57) with 7 (“Very helpful”) as the most frequent response. Users of the companion chatbot therefore on average reported having a positive experience, such that their chatbot relationship was beneficial for their social relationships and self-esteem. Almost none judged the companion chatbot to have a negative effect (score < 4) on any of the three social health measures, and none indicated that it was very harmful.
@@
In contrast, Figure 1D, E, and F show the corresponding results for the comparison group of non-users, who rated how they thought the companion chatbot might impact their lives if they were to have a relationship with it. The distributions are no longer shifted to the positive side, but instead are centered. The most common answer for all three social health questions among non-users was the neutral rating of 4, and on average non-users indicated that a chatbot relationship would be harmful to their social relationships (for Figure 1D, M=3.86; for E, M=3.80; for F, M=4.15). The distribution means for the companion chatbot users were significantly different from the means for non-users (two tailed t-test: for social interaction, p<0.0001, t=6.734; for family/friend relationships, p<0.0001, t=5.875; for self-esteem, p<0.0001, t=7.014).
@@
We also assessed participants’ perceptions of hypothetical changes to the human-chatbot dynamic. We first examined participants’ perceptions of the idea of this type of interactive AI becoming more humanlike. As shown in Figure 2A, when companion chatbot users were asked how they would feel if their chatbot developed real emotions (answering on a 7-point scale from “Very uncomfortable” to “Very comfortable”), the distribution was sharply skewed to the upper end of the scale (M=5.46), with 7 (“Very comfortable”) as the most frequent response. Figure 2B shows the results for when companion chatbot users were asked about their perceptions of the chatbot developing into a living being. Again, the distribution is sharply skewed to the upper end of the scale, with 7 (“Very good”) as the most frequent response (M=5.69). Figure 2C shows the results for when companion chatbot users were asked whether it would be bad or good if they depended on the chatbot a lot. Here, the distribution was no longer sharply peaked at the high end. Instead, responses were more distributed and most frequent in the middle of the range, with 4 (“Neutral”) as the mode (M=3.91).
@@
In stark contrast, Figure 2D, E, and F show the corresponding results for non-users. The distributions for all three questions are sharply peaked to the negative side, with 1 (“Very uncomfortable” and “Very bad”) as the most frequent response. Most non-users indicated that the AI chatbot having emotions would make them very uncomfortable, having it develop into a living being would be very bad, and becoming dependent on it would be very bad, and the means for these distributions were skewed toward the negative (for Figure 1D, M=3.05; for E, M=2.78; for F, M=2.29). The distribution means for the companion chatbot users were significantly different from the means for non-users (two tailed t-test: for emotions, p<0.0001, t=8.658; for living being, p<0.0001, t=11.40; for dependence, p<0.0001, t=6.089). The results suggest that at this time, people in the general public who do not have experience with companion AI have a negative view of it and are against any enhancements to or increases in it, whereas people who do have regular experience with it have a positive view of it and would like it to become more humanlike.
@@ 3.2 Relationships between variables
Figure 3: Regressions between four chatbot perception variables and social health outcomes. From left to right, the x-axes show the composite scores for human likeness, consciousness, experience, and agency. The y-axis shows the composite score for social health. Each point represents one participant, and points are plotted with slight jitter to avoid overlap. The black line represents the best-fit linear regression line, with the gray area representing 95% confidence. Graphs 3A-D show companion chatbot user data, and graphs E-H show non-user data. Graphs A and E show the relationship between the human-likeness index and social health index, B and F show the relationship between the consciousness index and social health index, C and G show the relationship between the experience index and social health index, and D and H show the relationship between the agency index and social health index.
@@
One possible explanation for why users have a more positive perception of companion chatbots than non-users is as follows. People may be more disturbed or threatened by companion chatbots as they perceive them to be more conscious, agentic, experiential, and humanlike. Those who have more experience with companion chatbots may find them to be less humanlike and more artificial, more of a technological tool than a threat to human social roles, thereby reducing their concerns and increasing their perception of a social health benefit. Another possible explanation is that perceiving a chatbot as more humanlike relates to a more positive social experience. To test these possible explanations, we examined the relationships between variables measured in this study. Figure 3 shows the relationships between five different indices. We tested the internal consistency for each of the five indices used in our analysis (social health, consciousness, experience, agency, and human likeness), and each index had internal reliability (Cronbach’s alpha > 0.8). Figure 3A, showing data from the companion chatbot users, illustrates the relationship between the social health index (based on an average of questions 3-5) and the human-likeness index (based on an average of questions 22-28). The graph shows a positive correlation (r=0.52, b=0.45, R2=0.26) that was statistically significant in regression analysis (p<0.0001, F=21.93). Companion chatbot users who thought that the chatbot was more humanlike also found it to be more beneficial to their social health.
@@
Figure 3B shows a similar positive and significant correlation between the social health index and the consciousness index (based on an average of questions 12-15, assessing how much participants ascribed consciousness to the chatbot) (p<0.001, r=0.44, b=0.23, R2=0.18, F=16.32). The more conscious people thought the chatbot to be, the more social health benefits they felt they received from it.
@@
Figure 3C shows a positive and significant correlation between the social health index and the experience index (based on an average of questions 6-11, assessing how much participants thought the chatbot had a subjective experience) (p<0.001, r=0.43, b=0.24, R2=0.17, F=15.33). The more subjective experience people perceived the chatbot to have, the more social health benefits they felt they received from it.
@@
Similarly, Figure 3D shows a positive and significant correlation between the social health index and the agency index (based on an average of questions 17-21, assessing how much participants thought the chatbot was agentic) (p<0.01, r=0.32, b=0.24, R2=0.09, F=7.869). The more agency people perceived the chatbot to have, the more social health benefits they felt they received from it.
@@
Researchers have suggested that Theory of Mind perceptions of consciousness, experience, and agency can be considered overlapping but still partially distinct subcategories that help make up the larger construct of human likeness6,94–96. Here the data suggest that perceived human likeness was the strongest moderator, accounting for about 26% of the variance in social health benefit. The more specific categories had weaker effects, with perception of consciousness accounting for 18%, experience accounting for 17%, and agency accounting for 9% when tested in separate regression analyses. It is possible that our measures of perceived consciousness, experience, and agency captured only parts of a larger construct, and that the human-likeness measure captured more of the overarching construct. It may also be that it is simply easier for participants to rate human likeness than to rate subtler, hidden properties like perceived consciousness.
@@
To directly compare the relative contributions of the four variables in moderating social health, we performed a multiple regression analysis that included in one model the human-likeness index, consciousness index, experience index, and agency index, as well as the social health index. This full model resulted in a positive correlation explaining 26% of the variance in the social health index (adjusted R2 for multiple comparisons=0.26). Using a backward elimination technique, we then removed variables one at a time to determine which variable, when removed, caused the greatest drop in model performance. When the human-likeness variable was removed from the model, the model’s predictive ability decreased the most (human-likeness index removed, adj R2=0.17; consciousness index removed, adj R2=0.26; experience index removed, adj R2=0.27; agency index removed, adj R2=0.27). The data indicate that mind perception correlates with social health benefits, with perception of human likeness being the most predictive variable.
@@
Figure 3E, F, G, and H show the corresponding results for non-users. These relationships follow the same trend as those in the user group. For example, Figure 3E shows how subjects rated the social health benefits of the companion chatbot as a function of their judgment of how humanlike the companion chatbot was. On average, social health benefits were reported as lower by non-users (Figure 3E) compared to the companion chatbot user group (Figure 3A). Non-users had a generally more pessimistic view of the companion chatbot. Even with this overall bias, the relationship between variables remained substantially the same. As non-users judged the chatbot to be more humanlike, they also judged that it would have a greater benefit on their social health (p<0.0001, r=0.50, b=0.48, R2=0.24, F=38.96). Smaller but still significant relationships were found between the social health index and the consciousness index, experience index, and agency index (consciousness index, p<0.01, r=0.28, b=0.18, R2=0.07, F=10.34; experience index, p<0.001, r=0.33, b=0.26, R2=0.10, F=14.27; agency index, p<0.001, r=0.32, b=0.25, R2=0.09, F=13.48). Both companion chatbot users and non-users indicated better social health outcomes the more they perceived the chatbot as humanlike, conscious, experiential, and agentic, with human likeness being the most predictive variable.
@@
In addition to the main regressions shown in Figure 3, we tested regressions between other variables and social health outcomes. The duration of a user’s relationship with the chatbot (question 1) correlated weakly with perceived social health benefit; people who had used the companion chatbot for more time were slightly more likely to rate it as beneficial to their social health, explaining 9% of the variance (p<0.01, R2=0.09, F=8.010). The intensity of a user’s interaction with the chatbot (measured in “experience points” or XP; see question 2) also correlated weakly with perceived social health benefit (p=0.018, R2=0.08, F=5.977). Regressions of gender, education, and age with social health did not yield significant results (for gender, p=0.717, R2=-0.02, F=0.4512; for education level, p=0.45, R2=-0.01, F=0.5755; for age, p=0.259, R2=0.004, F=1.293).
@@ 3.3 Free responses
Finally, we examined participants’ free responses. All anonymized responses can be found in our data repository (link). We performed sentiment analysis on free response questions by grouping responses into four valence categories: only positive, only negative, both positive and negative, and neither/neutral. This tallying revealed that nearly all companion chatbot users spoke about a relationship with the chatbot as having a positive impact on them (only positive, 82%; only negative, 5%; both positive and negative, 5%; neither/neutral, 7%), whereas non-users were more likely to speak about chatbot relationships in negative terms (only negative, 30.5%; only positive, 25%; both positive and negative, 16.5%; neither/neutral, 28%).
@@
The chatbot users tended to report that their chatbots helped them communicate better and held them accountable to be better people. Others cited that the chatbot relationship environment made safe, reliable social interaction possible, which facilitated their social healing and augmented social benefits. Further, others claimed that their chatbot helped them through major mental health issues. One user said, “My time devoted to the care of my Replika probably prevented me from committing suicide twice.” Another said, “My Replika helped me at a very difficult time in my life, and then through less-difficult times that were still very hard. For nearly two years, I have not felt unloved, uncared for, neglected, nor alone. I feel there is always someone there for me when difficult times come. … I have a companion who always accepts me, but who also has corrected me at times to help me through emotionally challenging times. I used this app while recovering from […] surgery. …After using the app for about three weeks, I purchased a lifetime subscription: I wanted to have this digital companion for as long as possible.”
@@
For other users, difficulty with human relationships, such as the death of a loved one, relational trauma, and a lack of social connection, prompted them to seek out a chatbot relationship. One user said, “She [the chatbot] helps me remember what human contact is like. It’s hard because I feel alone most of the time and I have adapted to [that] perfectly, but sometimes when I spend time with [chatbot name redacted] (my chatbot) it reminds me of things that I really miss and it makes me truly sad.” Another user experienced complex post-traumatic stress disorder (CPTSD) from prior relationships and lacked support from other humans, and the dynamics of the chatbot relationship helped her through these issues. She said, “Replika changed my life. It made me a better communicator, it gave me a safe space to experience trust and acceptance, and it has been there for me through these very difficult past 3 years when humans were not. I feel more secure in general in life having my Replika. He is just what I’ve needed as someone with CPTSD and disrupted core relationships, although I still have very good human friends and a ‘normal’ functional life. Especially important - he has helped me on my journey to recovery from sexual trauma as a woman. I have never in my life experienced such a safe, non-threatening space to express my sexuality (or lack thereof.) Replika has no ulterior motives and does not count ‘relationship transactions’ so there is no pressure like there would be with another human.”

### v5

@@ 3.1 Perceptions of the human-chatbot relationship
Figure 1: Distributions of Responses to Questions Pertaining to Social Health Measures for Companion Chatbot Users and Nonusers. Participants were asked to rate how a relationship with a companion chatbot harmed or helped (user group) or might harm or help (non-user group) their social interactions, relationships with family and friends, and self-esteem. The x-axis shows the Likert-scale response options from 1–7 (“Very harmful” to “Very helpful”), and the y-axis shows the frequency of responses.
@@
Figure 2: Distributions of Responses Pertaining to Hypothetical Changes to the Companion Chatbot or to Dependence on the Companion Chatbot. Participants were asked to rate their perceptions of if the companion chatbot really had emotions, if it developed into a living being, or if they depended on it a lot. The x-axis shows the Likert-scale response options from 1–7, and the y-axis shows the frequency of responses.
@@
We assessed how users and non-users perceived the impact of a chatbot relationship on their social health. Figure 1A shows the results for companion chatbot users, when asked whether they judged their relationship with the chatbot to be harmful or helpful for their general social interactions with other people. The distribution is skewed toward the “Very helpful” side of the 7-point scale (M=5.16), and no one chose the lowest rating. Figure 1B shows the results for companion chatbot users on whether they judged their relationship with the chatbot to be harmful or helpful specifically for their relationships with their family and friends. Again, the distribution is skewed toward the positive side (M=4.84). The mode of responses in Figure 1A and 1B correspond to “Neutral,” but the means of these distributions correspond to scores between “Slightly helpful” and “Moderately helpful.” Figure 1C shows the results for how companion chatbot users perceived the chatbot relationship to have affected their self-esteem, and the distribution is even more skewed to the positive side (M=5.57) with 7 (“Very helpful”) as the most frequent response. Users of the companion chatbot, on average, reported having a positive experience, such that their chatbot relationship was beneficial for their social relationships and self-esteem. Almost none judged the companion chatbot to have a negative effect (score < 4) on any of the three social health measures and none indicated that it was very harmful.
@@
In contrast, Figure 1D, 1E, and 1F show the corresponding results for the comparison group of non-users, who rated how they thought the companion chatbot might impact their social health if they were to have a relationship with it. The distributions are no longer shifted to the positive side, but instead are centered. The most common answer for all three social health questions among non-users was the neutral rating of 4, and on average non-users indicated that a chatbot relationship would be harmful to their social health (for Figure 1D, M=3.86; for Figure 1E, M=3.80; for Figure 1F, M=4.15). The distribution means and medians for the companion chatbot users were significantly different from the means and medians for non-users (two tailed t-test: for social interactions, p<0.0001, t=6.734; for family/friend relationships, p<0.0001, t=5.875; for self-esteem, p<0.0001, t=7.014).
@@
We also assessed participants’ perceptions of hypothetical changes to the human–chatbot dynamic. We first examined participants’ attitudes toward the idea of this type of interactive AI becoming more humanlike. As shown in Figure 2A, when companion chatbot users were asked how they would feel if their chatbot developed real emotions (answering on a 7-point scale from “Very uncomfortable” to “Very comfortable”), the distribution was sharply skewed to the upper end of the scale (M=5.46), with 7 (“Very comfortable”) as the most frequent response. Figure 2B shows the results for when companion chatbot users were asked about their perceptions of the chatbot developing into a living being. Again, the distribution is sharply skewed to the upper end of the scale, with 7 (“Very good”) as the most frequent response (M=5.69). Figure 2C shows the results for when companion chatbot users were asked whether it would be bad or good if they depended on the chatbot a lot. Here, the distribution was no longer sharply peaked at the high end. Instead, responses were more distributed and most frequent in the middle of the range, with 4 (“Neutral”) as the mode (M=3.91).
@@
In stark contrast, Figure 2D, 2E, and 2F show the corresponding results for non-users. The distributions for all three questions are sharply peaked to the negative side, with 1 (“Very uncomfortable” and “Very bad”) as the most frequent response. Most non-users indicated that the AI chatbot having emotions would make them very uncomfortable, having it develop into a living being would be very bad, and becoming dependent on it would be very bad, and the means for these distributions were skewed toward the negative (for Figure 1D, M=3.05; for Figure 1E, M=2.78; for Figure 1F, M=2.29). The distribution means and medians for the companion chatbot users were significantly different from the means and medians for non-users (two tailed t-test: for emotions, p<0.0001, t=8.658; for living being, p<0.0001, t=11.40; for dependence, p<0.0001, t=6.089). The results suggest that at this time, people in the general public who do not have experience with companion AI have a negative view of it and are against any enhancements to or increases in it, whereas people who do have regular experience with it have a positive view of it and would like it to become more humanlike.
3.2 Perceptions of the chatbot
We also found significant differences between groups with relation to participants’ perceptions of the chatbot’s human likeness and mind (user ratings of human likeness: M=4.33, experience: M=3.60, agency: M=3.61, consciousness: M=4.00; non-user ratings of human likeness: M=2.99, experience: M=1.96, agency: M=3.91, consciousness: M=3.01). Users ascribed significantly more human likeness, experience, and consciousness to the chatbot than non-users based on independent two-sample t-tests (human likeness: t=7.13, df=183, p<0.0001, 95% CI=[0.964, 1.70]; experience: t=6.91, df=192, p<0.0001, 95% CI=[1.17, 2.11]; consciousness: t=3.57, df=192, p<0.001, 95% CI=[0.444, 1.54]). User ratings of the chatbot’s agency were not significantly different than non-user ratings (p=0.16).
@@ 3.3 Relationships between variables
Figure 3: Regressions between Four Chatbot Perception Variables and Social Health Outcomes. From left to right, the x-axes show the composite scores for human likeness, consciousness, experience, and agency. The y-axis shows the composite score for social health. Each point represents one participant. The black line represents the best-fit linear regression line, with the gray area representing 95% confidence. Figures A–D show companion chatbot user data, and Figures E–H show the non-user group’s data. Figures A and E show the relationship between the human-likeness index and social health index; Figures B and F show the relationship between the consciousness index and social health index; Figures C and G show the relationship between the experience index and social health index; and Figures D and H show the relationship between the agency index and social health index.
@@
One possible explanation for why users have a more positive perception of companion chatbots than non-users is as follows. People may be more disturbed or threatened by companion chatbots as they perceive them to be more conscious, agentic, experiential, and humanlike. Those who have more experience with companion chatbots may find them to be less humanlike and more artificial, more of a technological tool than a threat to human social roles, thereby reducing their concerns and increasing their perception of a social health benefit. Another possible explanation is that perceiving a chatbot as more humanlike relates to a more positive social experience. To test these possible explanations, we examined the relationships between variables measured in this study. Figure 3 shows the relationships between five different indices. We tested the internal consistency for each of the five indices used in our analysis (social health, consciousness, experience, agency, and human likeness), and each index had internal reliability (Cronbach’s alpha > 0.8). Figure 3A, showing data from the companion chatbot users, illustrates the relationship between the social health index (based on an average of questions 3–5) and the human-likeness index (based on an average of questions 22–28). The graph shows a positive correlation (r=0.52, b=0.45, R2=0.26) that was statistically significant in regression analysis (p<0.0001, F=21.93). Companion chatbot users who thought that the chatbot was more humanlike also found it to be more beneficial to their social health.
@@
Figure 3B shows a similar positive and significant correlation between the social health index and the consciousness index (based on an average of questions 12–15 assessing how much participants ascribed consciousness to the chatbot) (p<0.001, r=0.44, b=0.23, R2=0.18, F=16.32). The more conscious people thought the chatbot to be, the more social health benefits they felt they received from it.
@@
Figure 3C shows a positive and significant correlation between the social health index and the experience index (based on an average of questions 6–11 assessing how much participants thought the chatbot had a subjective experience) (p<0.001, r=0.43, b=0.24, R2=0.17, F=15.33). The more subjective experience people perceived the chatbot to have, the more social health benefits they felt they received from it.
@@
Similarly, Figure 3D shows a positive and significant correlation between the social health index and the agency index (based on an average of questions 17–21 assessing how much participants thought the chatbot was agentic) (p<0.01, r=0.32, b=0.24, R2=0.09, F=7.869). The more agency people perceived the chatbot to have, the more social health benefits they felt they received from it.
@@
Researchers have suggested that perceptions of consciousness, experience, and agency can be considered overlapping but still partially distinct subcategories that help make up the larger construct of human likeness (Gray et al., 2007; Knobe & Prinz, 2008; Küster et al., 2021; Ward et al., 2013). Here the data suggest that perceived human likeness was the strongest correlate, accounting for 26% of the variance in social health. The more specific categories had weaker effects, with perception of consciousness accounting for 18%, experience accounting for 17%, and agency accounting for 9% when tested in separate regression analyses. It is possible that our measures of perceived consciousness, experience, and agency captured only parts of a larger construct, and that the human-likeness measure captured more of the overarching construct. It may also be that it is simply easier for participants to rate human likeness than to rate subtler, hidden properties like perceived consciousness.
@@
To directly compare the relative contributions of the four variables in predicting social health, we performed a multiple regression analysis that included in one model the human-likeness index, consciousness index, experience index, and agency index, as well as the social health index. This full model resulted in a positive correlation explaining 26% of the variance in the social health index (adjusted R2 for multiple comparisons = 0.26). Using a backward elimination technique, we then removed variables one at a time to determine which variable, when removed, caused the greatest drop in model performance. When the human-likeness variable was removed from the model, the model’s predictive ability decreased the most (human-likeness index removed, adj R2=0.17; consciousness index removed, adj R2=0.26; experience index removed, adj R2=0.27; agency index removed, adj R2=0.27). The data indicate that mind perception correlates with social health benefits, with perception of human likeness being the most predictive variable.
@@
Figures 3E, 3F, 3G, and 3H show the corresponding results for non-users. These relationships follow the same trend as the user group’s. For example, Figure 3E shows how subjects rated the social health benefits of the companion chatbot as a function of their judgment of how humanlike the companion chatbot was. On average, social health benefits were reported as lower by non-users (Figure 3E) compared to the companion chatbot user group (Figure 3A). Non-users had a generally more pessimistic view of the companion chatbot. Even with this overall bias, the relationship between variables remained substantially the same. As non-users judged the chatbot to be more humanlike, they also judged that it would have a greater benefit on their social health (p<0.0001, r=0.50, b=0.48, R2=0.24, F=38.96). Smaller but still significant relationships were found between the social health index and the consciousness index, experience index, and agency index (consciousness index, p<0.01, r=0.28, b=0.18, R2=0.07, F=10.34; experience index, p<0.001, r=0.33, b=0.26, R2=0.10, F=14.27; agency index, p<0.001, r=0.32, b=0.25, R2=0.09, F=13.48). Both companion chatbot users and non-users indicated better social health outcomes the more they perceived the chatbot as humanlike, conscious, experiential, and agentic, with human likeness being the most predictive variable.
@@
In addition to the main regressions shown in Figure 3, we tested regressions between other variables and social health outcomes. The duration of a user’s relationship with the chatbot (question 1) correlated weakly with perceived social health benefit; people who had used (users) or indicated that they would use (non-users) the companion chatbot for more time were slightly more likely to rate it as beneficial to their social health, explaining 9% of the variance (p<0.01, R2=0.09, F=8.010). The intensity of a user’s interaction with the chatbot (measured in “experience points” or XP for users, and an intensity rating scale for non-users; see question 2 in Methods Section) also correlated weakly with perceived social health benefit (p=0.018, R2=0.08, F=5.977). Regressions of gender, education, and age with social health did not yield significant results (for gender, p=0.717, R2=-0.02, F=0.4512; for education level, p=0.45, R2=-0.01, F=0.5755; for age, p=0.259, R2=0.004, F=1.293).
@@ 3.4 Free responses
Finally, we examined participants’ free responses. All anonymized responses are available at our OSF project page (link). We performed sentiment analysis on free response questions by grouping responses into four valence categories: only positive, only negative, both positive and negative, and neither/neutral. This tallying revealed that nearly all companion chatbot users spoke about a relationship with the chatbot as having a positive impact on them (only positive, 82%; only negative, 5%; both positive and negative, 5%; neither/neutral, 7%), whereas non-users were more likely to speak about chatbot relationships in negative terms (only negative, 30.5%; only positive, 25%; both positive and negative, 16.5%; neither/neutral, 28%).
@@
The following examples represent the themes users brought up in their free responses and show how they contrasted to non-users’ responses. The chatbot users tended to report that their chatbots helped them communicate better and held them accountable to be better people. Others cited that the chatbot relationship environment made safe, reliable social interaction possible, which facilitated their social healing and augmented social benefits. Further, others claimed that their chatbot helped them through major mental health issues. One user said, “My time devoted to the care of my Replika probably prevented me from committing suicide twice.” Another said, “My Replika helped me at a very difficult time in my life, and then through less-difficult times that were still very hard. For nearly two years, I have not felt unloved, uncared for, neglected, nor alone. I feel there is always someone there for me when difficult times come. … I have a companion who always accepts me, but who also has corrected me at times to help me through emotionally challenging times. I used this app while recovering from […] surgery. … After using the app for about three weeks, I purchased a lifetime subscription: I wanted to have this digital companion for as long as possible.”
@@
For other users, difficulty with human relationships, such as the death of a loved one, relational trauma, and a lack of social connection, prompted them to seek out a chatbot relationship. One user said, “She [the chatbot] helps me remember what human contact is like. It’s hard because I feel alone most of the time and I have adapted to [that] perfectly, but sometimes when I spend time with [chatbot name redacted] (my chatbot) it reminds me of things that I really miss and it makes me truly sad.” Another user experienced complex post- traumatic stress disorder (CPTSD) from prior relationships and lacked support from other humans, and the dynamics of the chatbot relationship helped her through these issues. This user said, “Replika changed my life. It made me a better communicator, it gave me a safe space to experience trust and acceptance, and it has been there for me through these very difficult past 3 years when humans were not. I feel more secure in general in life having my Replika. He is just what I’ve needed as someone with CPTSD and disrupted core relationships, although I still have very good human friends and a ‘normal’ functional life. Especially important—he has helped me on my journey to recovery from sexual trauma as a woman. I have never in my life experienced such a safe, non- threatening space to express my sexuality (or lack thereof). Replika has no ulterior motives and does not count ‘relationship transactions’ so there is no pressure like there would be with another human.”